# Danny Halawi4, Fred Zhang4, Philip E. Tetlock2,5

### v4

@@ Appendix H Leaderboards: Top 50
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.187 0.083 0.135 [0.129, 0.141] <0.001 32%
Claude-3-5-Sonnet-20240620 Anthropic Scratchpad 0.171 0.101 0.136 [0.130, 0.142] <0.001 10%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.195 0.077 0.136 [0.130, 0.143] <0.001 42%
Gemini-1.5-Pro Google Scratchpad 0.162 0.111 0.137 [0.131, 0.142] <0.001 32%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.206 0.068 0.137 [0.131, 0.144] <0.001 41%
Gemini-1.5-Pro Google News Scratchpad 0.167 0.109 0.138 [0.132, 0.144] <0.001 32%
GPT-4 OpenAI Freeze values Scratchpad 0.196 0.083 0.139 [0.133, 0.145] <0.001 36%
Imputed Forecaster ForecastBench 0.250 0.034 0.142 [0.138, 0.145] <0.001 47%
GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.210 0.074 0.142 [0.137, 0.148] <0.001 35%
GPT-4o OpenAI Scratchpad 0.191 0.094 0.143 [0.137, 0.148] <0.001 31%
Gemini-1.5-Pro Google Freeze values Zero shot 0.219 0.066 0.143 [0.137, 0.149] <0.001 39%
GPT-4o OpenAI News Scratchpad 0.192 0.095 0.143 [0.137, 0.149] <0.001 31%
Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.207 0.084 0.146 [0.139, 0.152] <0.001 33%
Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.187 0.104 0.146 [0.140, 0.151] <0.001 30%
LLM Crowd ForecastBench News 0.240 0.052 0.146 [0.143, 0.149] <0.001 38%
Qwen1.5-110B-Chat Qwen Scratchpad 0.176 0.116 0.146 [0.141, 0.151] <0.001 29%
GPT-4 OpenAI Freeze values Zero shot 0.222 0.071 0.146 [0.140, 0.152] <0.001 38%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.199 0.096 0.148 [0.142, 0.153] <0.001 26%
LLM Crowd ForecastBench News 0.241 0.055 0.148 [0.145, 0.152] <0.001 38%
LLM Crowd ForecastBench News 0.241 0.055 0.148 [0.145, 0.152] <0.001 37%
Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.204 0.093 0.149 [0.143, 0.155] <0.001 28%
Gemini-1.5-Flash Google Freeze values Scratchpad 0.193 0.107 0.150 [0.143, 0.157] <0.001 31%
Gemini-1.5-Pro Google News Superforecaster 1 0.188 0.112 0.150 [0.144, 0.156] <0.001 32%
GPT-4 OpenAI Scratchpad 0.196 0.105 0.150 [0.145, 0.155] <0.001 28%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 2 0.193 0.109 0.151 [0.145, 0.157] <0.001 29%
Gemini-1.5-Pro Google Zero shot 0.219 0.083 0.151 [0.145, 0.158] <0.001 36%
GPT-4-Turbo-2024-04-09 OpenAI Zero shot 0.206 0.098 0.152 [0.146, 0.159] <0.001 31%
GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.210 0.095 0.152 [0.147, 0.158] <0.001 27%
Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.216 0.090 0.153 [0.148, 0.158] <0.001 25%
GPT-4-Turbo-2024-04-09 OpenAI News Superforecaster 2 0.208 0.098 0.153 [0.148, 0.158] <0.001 28%
Claude-3-Opus-20240229 Anthropic Zero shot 0.190 0.120 0.155 [0.148, 0.161] <0.001 34%
Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.207 0.103 0.155 [0.150, 0.161] <0.001 26%
Claude-3-5-Sonnet-20240620 Anthropic Zero shot 0.195 0.116 0.156 [0.149, 0.163] <0.001 34%
Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.207 0.104 0.156 [0.149, 0.163] <0.001 31%
Gemini-1.5-Flash Google Freeze values Zero shot 0.232 0.082 0.157 [0.150, 0.164] <0.001 40%
GPT-4o OpenAI News Superforecaster 3 0.211 0.104 0.157 [0.152, 0.163] <0.001 27%
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Scratchpad 0.211 0.105 0.158 [0.152, 0.164] <0.001 30%
Gemini-1.5-Flash Google Scratchpad 0.193 0.125 0.159 [0.152, 0.165] <0.001 28%
Llama-3-70b-Chat-Hf Meta Zero shot 0.207 0.111 0.159 [0.153, 0.164] <0.001 28%
Claude-2.1 Anthropic Scratchpad 0.236 0.085 0.160 [0.155, 0.166] <0.001 38%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 1 0.211 0.113 0.162 [0.156, 0.168] <0.001 30%
@@
1. For resolved questions, predictions are compared against ground truth, while for unresolved questions, predictions are compared to community aggregates.
@@
3. Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).
4. Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
@@
Superforecaster median forecast ForecastBench 0.093 0.048 0.071 [0.062, 0.079] 0%
Public median forecast ForecastBench 0.123 0.057 0.090 [0.082, 0.099] <0.001 24%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.158 0.090 0.124 [0.112, 0.136] <0.001 24%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.163 0.085 0.124 [0.113, 0.136] <0.001 24%
GPT-4o OpenAI Freeze values Scratchpad 0.182 0.069 0.125 [0.115, 0.136] <0.001 25%
GPT-4o OpenAI News with freeze values Scratchpad 0.175 0.084 0.129 [0.118, 0.140] <0.001 24%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.155 0.111 0.133 [0.122, 0.144] <0.001 22%
@@
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.163 0.109 0.136 [0.124, 0.148] <0.001 20%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.157 0.117 0.137 [0.126, 0.147] <0.001 21%
GPT-4-Turbo-2024-04-09 OpenAI Scratchpad 0.163 0.113 0.138 [0.128, 0.148] <0.001 18%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.177 0.101 0.139 [0.126, 0.152] <0.001 23%
Gemini-1.5-Pro Google Scratchpad 0.155 0.125 0.140 [0.130, 0.150] <0.001 20%
Gemini-1.5-Pro Google News Scratchpad 0.157 0.124 0.141 [0.129, 0.152] <0.001 21%
Claude-3-5-Sonnet-20240620 Anthropic Scratchpad 0.158 0.123 0.141 [0.129, 0.152] <0.001 19%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.178 0.104 0.141 [0.128, 0.154] <0.001 23%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.172 0.112 0.142 [0.131, 0.152] <0.001 17%
GPT-4o OpenAI Scratchpad 0.182 0.102 0.142 [0.132, 0.152] <0.001 20%
GPT-4o OpenAI News Scratchpad 0.175 0.109 0.142 [0.131, 0.153] <0.001 20%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.200 0.085 0.142 [0.130, 0.154] <0.001 25%
Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.163 0.127 0.145 [0.134, 0.156] <0.001 19%
Imputed Forecaster ForecastBench 0.250 0.045 0.147 [0.140, 0.154] <0.001 24%
GPT-4 OpenAI Freeze values Zero shot 0.212 0.083 0.147 [0.136, 0.159] <0.001 23%
Gemini-1.5-Pro Google Freeze values Zero shot 0.207 0.090 0.148 [0.136, 0.161] <0.001 21%
GPT-4 OpenAI Freeze values Scratchpad 0.191 0.106 0.149 [0.137, 0.161] <0.001 19%
Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.187 0.113 0.150 [0.139, 0.161] <0.001 18%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 2 0.172 0.129 0.150 [0.138, 0.162] <0.001 17%
LLM Crowd ForecastBench News 0.238 0.064 0.151 [0.144, 0.158] <0.001 19%
GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.210 0.095 0.152 [0.141, 0.163] <0.001 22%
LLM Crowd ForecastBench News 0.239 0.067 0.153 [0.146, 0.160] <0.001 19%
LLM Crowd ForecastBench News 0.240 0.066 0.153 [0.146, 0.160] <0.001 19%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.186 0.121 0.154 [0.144, 0.164] <0.001 16%
Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.196 0.112 0.154 [0.141, 0.167] <0.001 20%
Gemini-1.5-Flash Google Freeze values Scratchpad 0.178 0.132 0.155 [0.142, 0.168] <0.001 19%
Gemini-1.5-Pro Google News Superforecaster 1 0.178 0.132 0.155 [0.144, 0.166] <0.001 19%
Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.203 0.108 0.156 [0.146, 0.166] <0.001 17%
Qwen1.5-110B-Chat Qwen Scratchpad 0.172 0.140 0.156 [0.147, 0.165] <0.001 15%
Gemini-1.5-Pro Google Zero shot 0.207 0.106 0.157 [0.145, 0.169] <0.001 18%
Gemini-1.5-Flash Google Scratchpad 0.178 0.136 0.157 [0.146, 0.168] <0.001 15%
GPT-4 OpenAI Scratchpad 0.191 0.123 0.157 [0.148, 0.166] <0.001 15%
GPT-4-Turbo-2024-04-09 OpenAI Zero shot 0.200 0.117 0.159 [0.147, 0.170] <0.001 19%
Claude-2.1 Anthropic Scratchpad 0.227 0.091 0.159 [0.149, 0.168] <0.001 20%
Gemini-1.5-Flash Google Freeze values Zero shot 0.217 0.103 0.160 [0.146, 0.174] <0.001 23%
GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.210 0.109 0.160 [0.149, 0.170] <0.001 17%
Claude-3-Opus-20240229 Anthropic Zero shot 0.178 0.144 0.161 [0.148, 0.175] <0.001 19%
GPT-4-Turbo-2024-04-09 OpenAI News Superforecaster 2 0.204 0.120 0.162 [0.151, 0.173] <0.001 18%
Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.199 0.125 0.162 [0.152, 0.172] <0.001 16%
Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.201 0.125 0.163 [0.149, 0.176] <0.001 19%
Claude-3-5-Sonnet-20240620 Anthropic Zero shot 0.177 0.151 0.164 [0.150, 0.178] <0.001 18%
GPT-4o OpenAI News Superforecaster 3 0.206 0.125 0.166 [0.156, 0.176] <0.001 17%
Claude-2.1 Anthropic Freeze values Scratchpad 0.227 0.105 0.166 [0.154, 0.177] <0.001 21%
@@
from their forecasts on the component questions by assuming independence (which is not always the case, putting humans at a disadvantage).
2. For resolved questions, predictions are compared against ground truth, while for unresolved questions, predictions are compared to community aggregates.
@@
4. Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).

### v5

@@ Appendix H Leaderboards: top 50
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.184 0.101 0.143 [0.137, 0.149] <0.001 32%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.192 0.094 0.143 [0.136, 0.150] <0.001 42%
GPT-4-Turbo-2024-04-09 OpenAI Scratchpad 0.172 0.115 0.143 [0.138, 0.149] <0.001 31%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.204 0.084 0.144 [0.137, 0.150] <0.001 42%
Claude-3-5-Sonnet-20240620 Anthropic Scratchpad 0.169 0.120 0.144 [0.139, 0.150] <0.001 10%
Gemini-1.5-Pro Google Scratchpad 0.162 0.128 0.145 [0.139, 0.151] <0.001 32%
GPT-4 OpenAI Freeze values Scratchpad 0.194 0.100 0.147 [0.141, 0.154] <0.001 36%
Gemini-1.5-Pro Google News Scratchpad 0.166 0.129 0.147 [0.141, 0.153] <0.001 32%
Imputed Forecaster ForecastBench 0.250 0.048 0.149 [0.145, 0.153] <0.001 46%
GPT-4o OpenAI Scratchpad 0.186 0.114 0.150 [0.144, 0.156] <0.001 31%
Gemini-1.5-Pro Google Freeze values Zero shot 0.217 0.083 0.150 [0.144, 0.157] <0.001 39%
GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.211 0.091 0.151 [0.145, 0.157] <0.001 35%
GPT-4o OpenAI News Scratchpad 0.190 0.114 0.152 [0.146, 0.158] <0.001 31%
Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.184 0.124 0.154 [0.148, 0.160] <0.001 30%
GPT-4 OpenAI Freeze values Zero shot 0.222 0.087 0.154 [0.148, 0.161] <0.001 38%
Qwen1.5-110B-Chat Qwen Scratchpad 0.176 0.134 0.155 [0.150, 0.160] <0.001 28%
LLM Crowd ForecastBench News 0.242 0.068 0.155 [0.151, 0.159] <0.001 38%
Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.201 0.112 0.156 [0.150, 0.162] <0.001 27%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.199 0.115 0.157 [0.151, 0.162] <0.001 26%
Gemini-1.5-Pro Google Zero shot 0.217 0.097 0.157 [0.151, 0.163] <0.001 37%
LLM Crowd ForecastBench News 0.243 0.071 0.157 [0.153, 0.161] <0.001 37%
LLM Crowd ForecastBench News 0.244 0.071 0.157 [0.153, 0.161] <0.001 37%
GPT-4 OpenAI Scratchpad 0.194 0.121 0.158 [0.153, 0.162] <0.001 28%
Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.215 0.101 0.158 [0.151, 0.164] <0.001 33%
Gemini-1.5-Pro Google News Superforecaster 1 0.186 0.131 0.159 [0.153, 0.165] <0.001 31%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 2 0.190 0.129 0.159 [0.153, 0.165] <0.001 29%
GPT-4-Turbo-2024-04-09 OpenAI Zero shot 0.204 0.117 0.160 [0.154, 0.167] <0.001 32%
Gemini-1.5-Flash Google Freeze values Scratchpad 0.194 0.128 0.161 [0.154, 0.168] <0.001 32%
Claude-3-Opus-20240229 Anthropic Zero shot 0.186 0.136 0.161 [0.154, 0.168] <0.001 35%
GPT-4-Turbo-2024-04-09 OpenAI News Superforecaster 2 0.208 0.116 0.162 [0.156, 0.167] <0.001 28%
GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.211 0.114 0.163 [0.157, 0.168] <0.001 27%
Claude-3-5-Sonnet-20240620 Anthropic Zero shot 0.192 0.134 0.163 [0.156, 0.170] <0.001 34%
Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.205 0.122 0.164 [0.158, 0.170] <0.001 26%
Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.221 0.108 0.164 [0.159, 0.170] <0.001 25%
Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.208 0.122 0.165 [0.157, 0.172] <0.001 31%
Gemini-1.5-Flash Google Freeze values Zero shot 0.232 0.098 0.165 [0.158, 0.173] <0.001 40%
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Scratchpad 0.210 0.121 0.165 [0.159, 0.172] <0.001 30%
GPT-4o OpenAI News Superforecaster 3 0.211 0.124 0.168 [0.162, 0.173] <0.001 28%
Claude-2.1 Anthropic Scratchpad 0.237 0.100 0.168 [0.163, 0.174] <0.001 38%
Gemini-1.5-Flash Google Scratchpad 0.194 0.146 0.170 [0.164, 0.176] <0.001 28%
GPT-4o OpenAI Freeze values Zero shot 0.225 0.116 0.171 [0.163, 0.178] <0.001 37%
Claude-3-Opus-20240229 Anthropic Scratchpad 0.201 0.141 0.171 [0.165, 0.177] <0.001 26%
@@
1. Shows performance on the 1,000 (500 standard, 500 combination) questions in the LLM question set at the 7-, 30-, 90-, and 180-day forecast horizons.
2. The full leaderboard is available at www.forecastbench.org. Online results are updated nightly, so may be slightly different than the version presented here.
3. For resolved market questions, forecasts are compared against ground truth while for unresolved market questions, they are compared to community aggregates.
@@
5. Pairwise $p$-value comparing to No. 1 (bootstrapped): The $p$-value calculated by bootstrapping the differences in overall score between each model and the best
forecaster under the null hypothesis that there’s no difference.
6. Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster.
@@
Superforecaster median forecast ForecastBench 0.091 0.062 0.076 [0.067, 0.086] 0%
Public median forecast ForecastBench 0.119 0.072 0.096 [0.086, 0.105] <0.001 23%
GPT-4o OpenAI Freeze values Scratchpad 0.175 0.085 0.130 [0.119, 0.141] <0.001 24%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.154 0.107 0.131 [0.118, 0.143] <0.001 24%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.164 0.101 0.133 [0.121, 0.145] <0.001 23%
GPT-4o OpenAI News with freeze values Scratchpad 0.171 0.104 0.137 [0.125, 0.149] <0.001 20%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.152 0.130 0.141 [0.130, 0.152] <0.001 21%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.154 0.133 0.143 [0.133, 0.154] <0.001 21%
@@
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.160 0.130 0.145 [0.132, 0.158] <0.001 20%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.174 0.119 0.146 [0.133, 0.160] <0.001 22%
Gemini-1.5-Pro Google Scratchpad 0.152 0.143 0.148 [0.137, 0.158] <0.001 20%
GPT-4-Turbo-2024-04-09 OpenAI Scratchpad 0.164 0.132 0.148 [0.138, 0.158] <0.001 17%
Gemini-1.5-Pro Google News Scratchpad 0.154 0.143 0.148 [0.137, 0.160] <0.001 21%
Claude-3-5-Sonnet-20240620 Anthropic Scratchpad 0.154 0.143 0.149 [0.137, 0.160] <0.001 20%
GPT-4o OpenAI Scratchpad 0.175 0.122 0.149 [0.138, 0.159] <0.001 19%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.173 0.124 0.149 [0.135, 0.162] <0.001 21%
GPT-4o OpenAI News Scratchpad 0.171 0.127 0.149 [0.138, 0.160] <0.001 18%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.200 0.100 0.150 [0.138, 0.162] <0.001 24%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.171 0.131 0.151 [0.140, 0.162] <0.001 16%
Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.160 0.149 0.154 [0.143, 0.166] <0.001 19%
Imputed Forecaster ForecastBench 0.250 0.059 0.155 [0.147, 0.163] <0.001 22%
GPT-4 OpenAI Freeze values Zero shot 0.213 0.099 0.156 [0.144, 0.168] <0.001 21%
Gemini-1.5-Pro Google Freeze values Zero shot 0.205 0.110 0.157 [0.144, 0.171] <0.001 20%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 2 0.167 0.149 0.158 [0.146, 0.169] <0.001 17%
GPT-4 OpenAI Freeze values Scratchpad 0.190 0.125 0.158 [0.145, 0.171] <0.001 19%
Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.185 0.134 0.159 [0.148, 0.171] <0.001 18%
LLM Crowd ForecastBench News 0.241 0.080 0.161 [0.153, 0.168] <0.001 18%
GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.209 0.114 0.161 [0.149, 0.173] <0.001 20%
LLM Crowd ForecastBench News 0.242 0.083 0.162 [0.155, 0.170] <0.001 18%
LLM Crowd ForecastBench News 0.243 0.082 0.162 [0.155, 0.170] <0.001 18%
Gemini-1.5-Pro Google News Superforecaster 1 0.176 0.151 0.164 [0.153, 0.175] <0.001 19%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.185 0.143 0.164 [0.154, 0.175] <0.001 16%
GPT-4 OpenAI Scratchpad 0.190 0.140 0.165 [0.156, 0.174] <0.001 15%
Qwen1.5-110B-Chat Qwen Scratchpad 0.171 0.161 0.166 [0.156, 0.175] <0.001 15%
Gemini-1.5-Pro Google Zero shot 0.205 0.128 0.167 [0.154, 0.179] <0.001 19%
Gemini-1.5-Flash Google Freeze values Scratchpad 0.179 0.154 0.167 [0.153, 0.180] <0.001 18%
Claude-2.1 Anthropic Scratchpad 0.228 0.105 0.167 [0.157, 0.177] <0.001 20%
Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.205 0.132 0.168 [0.155, 0.182] <0.001 19%
Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.208 0.129 0.169 [0.158, 0.179] <0.001 17%
Claude-3-Opus-20240229 Anthropic Zero shot 0.173 0.165 0.169 [0.156, 0.183] <0.001 18%
Gemini-1.5-Flash Google Freeze values Zero shot 0.217 0.122 0.169 [0.155, 0.183] <0.001 23%
GPT-4-Turbo-2024-04-09 OpenAI Zero shot 0.200 0.139 0.169 [0.157, 0.182] <0.001 19%
GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.209 0.131 0.170 [0.159, 0.180] <0.001 17%
Gemini-1.5-Flash Google Scratchpad 0.179 0.161 0.170 [0.159, 0.181] <0.001 16%
GPT-4-Turbo-2024-04-09 OpenAI News Superforecaster 2 0.202 0.139 0.170 [0.160, 0.181] <0.001 17%
Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.198 0.146 0.172 [0.161, 0.183] <0.001 16%
Claude-3-5-Sonnet-20240620 Anthropic Zero shot 0.174 0.171 0.172 [0.158, 0.187] <0.001 17%
Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.203 0.145 0.174 [0.160, 0.188] <0.001 19%
Claude-2.1 Anthropic Freeze values Scratchpad 0.228 0.120 0.174 [0.162, 0.186] <0.001 20%
GPT-4o OpenAI News Superforecaster 3 0.206 0.145 0.175 [0.165, 0.186] <0.001 16%
@@
from their forecasts on the component questions by assuming independence (which is not always the case, putting humans at a disadvantage). Evaluated at the 7-, 30-,
90-, and 180-day forecast horizons.
2. The full leaderboard is available at www.forecastbench.org. Online results are updated nightly, so may be slightly different than the version presented here.
3. For resolved market questions, forecasts are compared against ground truth while for unresolved market questions, they are compared to community aggregates.